a rectangular table is 4 feet longer as long as it is wide. If the area is 144 ft to the 2nd power find the length and width.
Found 2 solutions by solver91311, mananth:
Answer by solver91311(24713)



The area of a rectangle is length times width. If you let represent the width, then has to represent the length. Therefore:





Just solve the quadratic. You will need to use the quadratic formula. Discard the negative root because you are looking for a positive measure of length.

John


Answer by mananth(16949)

let the width be x feet
length = x+4 feet
Area = L*W
..
x(x+4)=144
x^2+4x-144=0
..
Find X1, X2 the roots
X1= -4 +sqrt(16+576) / 2
= 10.165
X2 =-4 -sqrt(16+576) / 2
= -14.165
..
X= 10.165 feet the width
Length = 14.165 feet
.
CHECK
14.165*10.165== 143.98